# LocDiffusion: Identifying Locations on Earth by Diffusing in the Hilbert Space

### v1

@@ 4.1 Problem Setup and Intuitions
As we have outlined in the introduction, our goal is to find a position encoding method that encodes the spherical surface into a dense subset of $\mathbb{R}^{d}$ (ideally the entire $\mathbb{R}^{d}$) and accurately decode points back to spherical coordinates. There are several mathematical properties such position encoding and decoding method should have. For rigorous discussions, we give definitions of the aforementioned properties and demonstrate how they guide the finding of our SHDD encoding-decoding framework.
@@
Since we are projecting a set of 2-dimensional points in $\mathcal{C}$ into a high-dimensional Euclidean space $\mathcal{S}_{\mathbb{PE}}$, dense filling is impossible. However, if we define a difference measure $\mathcal{E}:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}$, then $\mathcal{S}_{\mathbb{PE}}$ can be partitioned by the following equivalence relation:
@@
then the sparsity problem is resolved, since now diffusion in $\mathcal{S}_{\mathbb{PE}}$ equals a random walk among spherical points and small perturbation will not result in an abrupt jump on the spherical surface.
@@
It is not an easy task to find such $\mathcal{E}$, especially considering computational constraints (e.g., it is impossible to exactly evaluate the $\operatorname*{arg\,min}$ function in Equation 3). Fortunately, we find that by treating spherical points as special spherical functions and represent them using Spherical Harmonics coefficients, we can define $\mathcal{E}$ as spherical KL-divergence which satisfies all the desirable properties mentioned above, thus addressing the sparsity and the non-linearity problems as a whole. Moreover, the choice of Spherical Harmonics coefficients also enables efficient computation.
@@ 4.2 Spherical Harmonics Dirac Delta (SHDD) Encoding
As discussed in Section 3.2, we can represent spherical points as spherical Dirac delta functions. Consider Section 3.1, a spherical Dirac delta function can be encoded as an infinite-dimensional real-valued coefficient vector, i.e. a point in a Hilbert space. $\bigcup$ denotes vector concatenation.
$\mathbb{PE}_{\text{SHDD}}(\theta_{0},\phi_{0}):=\bigcup_{l=0}^{\infty}\bigcup_{m=-l}^{l}\left[C_{lm}\right]$ (7)
Thus, the spherical harmonics coefficient vector can be used to uniquely represent a point $(\theta_{0},\phi_{0})$ on the sphere. In practice, it is impossible to deal with infinite-dimensional coefficient vectors. It is also impossible to deal with the infinite probability density of spherical Dirac delta functions. These two practical constraints, fortunately, can be settled as a whole: we truncate the coefficient vector up to its leading $(L+1)^{2}$ dimensions, where $L$ is the maximum degree of associate Legendre polynomials. Therefore, the $L$-degree representation of point $(\theta_{0},\phi_{0})$ is defined as
@@
We call this $(L+1)^{2}$-dimensional real-valued vector the $L$-degree Spherical Harmonics Dirac Delta (SHDD) Representation of $(\theta_{0},\phi_{0})$ and $\mathbb{PE}_{\text{SHDD}}$ the SHDD encoder. Each SHDD representation corresponds to an approximation of the true spherical Dirac delta function $\delta_{(\theta_{0},\phi_{0})}$, whose probability density concentrates in a region surrounding $(\theta_{0},\phi_{0})$ rather than a single point, solving the infinite density problem. The Legendre polynomials have finer granularity as their degree $L$ increases, which makes SHDD representations, like other frequency-based location encoding methods such as Sphere2Vec (Mai et al., 2023b), capable of capturing multi-scale spatial information.
Problems remain on how to find values of $C_{lm}$. For arbitrary spherical functions, $C_{lm}$ needs to be iteratively computed. However, for spherical Dirac delta functions, we can efficiently obtain $C_{lm}$ given the fact that the coefficients of the Legendre polynomials are the values of the Legendre polynomials at $(\theta_{0},\phi_{0})$ (Arfken et al., 2011), i.e.,
$F$ $\equiv\delta_{(\theta_{0},\phi_{0})}\Leftrightarrow\forall(\theta,\phi),$ (9)
$F(\theta,\phi)$ $=\sum_{l=0}^{\infty}\sum_{m=-l}^{l}Y_{lm}(\theta_{0},\phi_{0})Y_{lm}(\theta,\phi)$ (10)
That is, for spherical Dirac delta functions, $C_{lm}\equiv Y_{lm}(\theta_{0},\phi_{0})$ for any $l$ and $m$. So instead of iteratively computing $\{C_{lm}\}$ in the general case, the encoding procedure can be reduced to a simple look-up of $Y_{lm}$ values.
It is worth noting that while the SHDD representation in Equation 8 has the same expression as the SH positional encoding used in a recent work (Rußwurm et al., 2024), they refer to distinct mathematical objects. The SH positional encoding of a point $(\theta_{0},\phi_{0})$ in their work is the sequence of evaluated $Y_{lm}(\theta_{0},\phi_{0})$ values, which forms a sparse feature space. The SHDD representation of a point $(\theta_{0},\phi_{0})$ in our work is the $C_{lm}$ values of the corresponding spherical Dirac delta function $\delta_{(\theta_{0},\phi_{0})}$, whose $\mathcal{E}$-equivalence classes form a Hilbert coefficient space. The reason that the two types of positional encodings coincidentally have identical expressions is only because spherical Dirac functions satisfy Equation 9, i.e., $C_{lm}\equiv Y_{lm}(\theta_{0},\phi_{0})$.
4.3 The SHDD Distance Measure
SHDD KL-Divergence   An $L$-degree SHDD representation corresponds to a spherical Dirac delta function $\delta$, and an arbitrary $\mathbb{R}^{(L+1)^{2}}$ vector corresponds to certain spherical function $F$. Thus, we can use the reverse KL-divergence between (the normalized) $F$ and $\delta$ as the difference measure $\mathcal{E}$. Let $p_{(\theta,\phi)}$ and $q_{\mathbf{e}}$ be the normalized probability distributions corresponding to the SHDD representation of $(\theta,\phi)$ and an arbitrary $\mathbb{R}^{(L+1)^{2}}$ vector $\mathbf{e}=\bigcup_{l=0}^{L}\bigcup_{m=-l}^{l}\left[e_{lm}\right]$
$p_{(\theta,\phi)}(u,v):=$ $\exp\left(\sum_{l=0}^{L}\sum_{m=-l}^{l}Y_{lm}(\theta,\phi)Y_{lm}(u,v)\right)$
$/Z(\mathbb{PE}_{\text{SHDD}}(\theta,\phi))$ (11)
@@ 4.3 The SHDD Distance Measure
Here
$Z(\mathbf{e})=\int_{0}^{\pi}\int_{0}^{2\pi}\exp\left(\sum_{l=0}^{L}\sum_{m=-l}^{l}e_{lm}Y_{lm}(u^{{}^{\prime}},v^{{}^{\prime}})\right)\mathbf{d}u^{{}^{\prime}}\mathbf{d}v^{{}^{\prime}}$
is a normalization constant and the exponential ensures that probabilities are non-negative. The SHDD KL-divergence
$\mathcal{L}_{\text{SHDD-KL}}$ $(\mathbf{e},\mathbb{PE}_{\text{SHDD}}((\theta,\phi))):=$ (13)
$\int_{0}^{\pi}\int_{0}^{2\pi}q_{\mathbf{e}}(u,v)\log\dfrac{q_{\mathbf{e}}(u,v)}{p_{(\theta,\phi)}(u,v)}\mathbf{d}u\mathbf{d}v$
It is easy to verify that the SHDD KL-divergence is a continuous difference measure. As for the property described in Equation 6, notice that by (Gibbs & Su, 2002), the Wasserstein-2 distance $W_{2}$ between $p_{(\theta,\phi)}$ and $q_{\mathbf{e}}$ is bounded by the KL-divergence in the following inequality:
@@
$C$ being a finite constant. $W_{2}$, being the Earth Mover’s Distance, quantifies the amount of probability mass transport between two distributions. Thus, when $\mathcal{L}_{\text{SHDD-KL}}(\mathbf{e},\mathbb{PE}_{\text{SHDD}}((\theta,\phi))$ is small, the difference in probability mass distribution is also small, and consequently the largest-mass-region found by the mode-seeking SHDD decoder will also remain mostly unchanged. Figure 1 visualizes this with concrete examples (pretrained Sphere2Vec location encoder and learned neural decoder v.s. our SHDD encoder and decoder) using heatmaps.
@@ 4.4 SHDD Decoding
It is impractical to compute $\mathbb{PD}_{\text{KL}}$ exactly. Luckily, Equation 9 makes a natural simplification possible. Notice that minimizing reverse KL-divergence leads to mode-seeking behavior (Minka et al., 2005), i.e. the $(\theta,\phi)$ that satisfies Equation 15 should fall within the region with the largest probability mass. Thus, we can decode $\mathbf{e}$ by finding the center of its probability mass concentration.
@@
$\mathbb{PD}_{\text{mode}}$ $(\mathbf{e};\rho):=\operatorname*{arg\,max}_{(\theta,\phi)}\bigg{\{}\int_{\theta-\rho}^{\theta+\rho}\int_{\phi-\rho}^{\phi+\rho}$
$\exp\left(\sum_{l=0}^{L}\sum_{m=-l}^{l}e_{lm}Y_{lm}(u,v)\right)\mathbf{d}u\mathbf{d}v\bigg{\}}.$ (16)
@@
One advantage of adopting the SHDD decoder is its learning-free property. Unlike learned neural decoders, there is no loss introduced during the decoding stage. Besides, the mapping from diffusion outputs to spherical coordinates is shown to be continuous and relatively smooth. Therefore, it is safe to train latent diffusion models only using the SHDD KL-divergence loss $\mathcal{L}_{\text{SHDD-KL}}$.
Another critical advantage is that the spatial resolution of our SHDD decoder is arbitrary (i.e., real-valued), and not dependent on partitions of the spherical surface or the spatial distributions of image/location galleries. This is because the SHDD representation is in effect a continuous spherical function and in theory one can evaluate it in arbitrary resolution. The only two constraints are the maximum degree of Legendre polynomials $L$ which limits the spatial resolution of the spherical function itself and the computational resources (e.g., float32 or float64, evaluation granularity $\rho$), both being independent from other factors.
@@ 4.5 Conditional SirenNet-Based UNet (CS-UNet)
Figure 2: (a): The architecture of Condition SirenNet Module (C-Siren). $x$ is the input latent vector, $x^{{}^{\prime}}$ is the output latent vector, $t$ is the scalar timestep, and $e_{I}$ is the embedding of the input image. $d_{i}$ is the input dimension, $d_{o}$ is the output dimension, $d_{T}$ is the time embedding dimension, $d_{I}$ is the conditional embedding dimension. (b): The architecture of Conditional SirenNet-Based UNet (CS-UNet) and the workflow of LocDiffusion. $d$ is the latent dimension. The numbered circles denote the order of training steps.
@@
Inspired by (Rußwurm et al., 2024), we used SirenNet (Sitzmann et al., 2020) as the backbone of our diffusion model. The theoretical motivation behind this decision is that Spherical Harmonics coefficients are sums of sinusoidal and cosinusoidal functions (See Appendix A.2). Using sine as the activation function helps preserve gradients because the derivatives of sinusoidal/cosinusoidal functions are still sinusoidal and cosinusoidal functions. Figure 2 depicts the network architecture of the Conditional SirenNet (C-Siren) module. The design is straightforward: inputs are the latent vector $x$, the image condition embedding $e_{I}$, and the diffusion step $t$. First, we use feed-forward layers to project $x$ and $e_{I}$ into hidden vectors $h_{x}$, $h_{I}$. Then we use the sinusoidal embedding layer (Song et al., 2021) and feed-forward layers to project the discrete diffusion timestep $t$ into a scale vector $\alpha_{t}$ and a shift vector $\epsilon_{t}$. Then, we transform $h_{x}$ into $h_{x}=(1+\alpha_{t})\odot h_{x}+\epsilon_{t}$, which is an unconditional denoising step. Following that, we sum the transformed $h_{x}$ and the condition $h_{I}$ and pass the sum to a feed-forward layer, which adjusts denoising step under the guidance of the condition. Finally, output the sine-activated hidden vector to the next C-Siren module. Figure 2 describes the architecture of the Conditional SirenNet-Based Unet.
4.6 LocDiffusion
@@
Next, we introduce the training cycle of our LocDiffusion model as illustrated in Figure 2. A training data sample includes an input image $\mathbf{I}$ and its associated geolocation $\mathbf{p}=(\theta,\phi)$ serving as the prediction target. First, we use a frozen CLIP-based image encoder (Radford et al., 2021) to encode the image $\mathbf{I}$ into an image embedding $e_{I}$. Then, we encode the geolocation $\mathbf{p}=(\theta,\phi)$ into its SHDD representation $\mathbb{PE}_{\text{SHDD}}(\theta,\phi)$ and store them in a look-up table. Following that, we perform a standard DDPM training (Ho et al., 2020a) based on the proposed CS-UNet architecture as shown in Figure 2. In a forward pass in the latent diffusion process, the spherical Dirac delta function $\delta_{(\theta_{0},\phi_{0})}$ defined by $\mathbb{PE}_{\text{SHDD}}(\theta,\phi)$ will be gradually added noise until being reduced to a vector whose values in each dimension are purely generated from Gaussian noise. In a backward pass of the latent diffusion model, the CS-UNet will start with a noise vector and gradually recover the $\delta_{(\theta_{0},\phi_{0})}$ spherical function. We implement the DDPM algorithm based on the open-source PyTorch implementation. We use the SHDD KL-divergence $\mathcal{L}_{\text{SHDD-KL}}$ between the ground-truth SHDD representation $\mathbb{PE}_{\text{SHDD}}(\theta,\phi)$ and the diffusion output as the training objective, because it is more computationally stable and preserves the spatial multi-scalability than the spherical MSE (e.g. great circle distance) loss. During inferencing, we sample coefficient vectors from Gaussian noise conditioned on CLIP-based image embeddings and use $\mathbb{PD}_{\text{mode}}$ to predict locations.
There are two important implementation details worth mentioning. In practice, the integrals in Equation 13 and Equation 4.4 are approximated by summation. More specifically, we select a set of $N$ anchor points $\mathcal{A}_{N}=\{(\theta_{i},\phi_{i})\in\mathcal{C}\}_{i=1}^{N}$ on the sphere, and
$\mathcal{L}_{\text{SHDD-KL}}$ $(\mathbf{e},\mathbb{PE}_{\text{SHDD}}(\theta,\phi))=$ (17)
$\sum_{i=1}^{N}q_{\mathbf{e}}(\theta_{i},\phi_{i})\times\log\dfrac{q_{\mathbf{e}}(\theta_{i},\phi_{i})}{p_{(\theta,\phi)}(\theta_{i},\phi_{i})}.$
@@
$\mathbb{PD}_{\text{mode}}(\mathbf{e};\rho)=$ $\operatorname*{arg\,max}_{(\theta,\phi)}\sum_{i=1}^{N}\mathbb{I}\{d_{\mathcal{C}}((\theta,\phi),(\theta_{i},\phi_{i}))<\rho\}$
$\times$ $\exp\left(\sum_{l=0}^{L}\sum_{m=-l}^{l}e_{lm}Y_{lm}(\theta_{i},\phi_{i})\right).$ (18)
@@
$\mathcal{L}_{\text{SHDD-KL}}$ is used for training, thus we random sample $N=2048$ anchor points over the globe for each mini-batch to avoid overfitting. As for $\mathbb{PD}_{\text{mode}}$, the choice of $\mathcal{A}_{N}$ introduces inductive bias – the regions with more anchor points have heavier impact on the decoding results and higher spatial resolutions. However, Table 2 shows that LocDiffusion performs stably well on different $\mathcal{A}_{N}$.

### v2

@@ 4.1 Problem Setup and Intuitions
As we have outlined in the introduction, our goal is to find a position encoding method that encodes the spherical surface into a dense subset of $\mathbb{R}^{d}$ (ideally the entire $\mathbb{R}^{d}$) and accurately decodes points back to spherical coordinates. There are several mathematical properties such position encoding and decoding method should have. For rigorous discussions, we give definitions of the aforementioned properties and demonstrate how they guide the finding of our SHDD encoding-decoding framework.
@@
Since we are projecting a set of 2-dimensional points in $\mathcal{C}$ into a high-dimensional Euclidean space $\mathcal{S}_{\mathbb{PE}}$, dense filling is impossible.
However, if we define a difference measure $\mathcal{E}:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}$, then $\mathcal{S}_{\mathbb{PE}}$ can be partitioned by the following equivalence relation:
@@
then the sparsity problem is resolved. Since now diffusion in $\mathcal{S}_{\mathbb{PE}}$ equals a random walk among spherical points and a small perturbation in $\mathcal{S}_{\mathbb{PE}}$ will not result in an abrupt jump on $\mathcal{C}$.
@@
It is not an easy task to find such $\mathcal{E}$, especially considering computational constraints (e.g., it is impossible to exactly evaluate the $\operatorname*{arg\,min}$ function in Equation 1). Fortunately, we find that by treating spherical points as special spherical functions and representing them using Spherical Harmonics coefficients, we can define $\mathcal{E}$ as spherical KL-divergence which satisfies all the desirable properties mentioned above, thus addressing the sparsity and the non-linearity problems as a whole. Moreover, the choice of Spherical Harmonics coefficients also enables efficient computation.
@@ 4.2 Spherical Harmonics Dirac Delta (SHDD) Encoding
As discussed in Section 3, we can represent spherical points as spherical Dirac delta functions, and a spherical Dirac delta function can be encoded as an infinite-dimensional real-valued coefficient vector, i.e. a point in a Hilbert space: $\mathbb{PE}_{\text{SHDD}}(\theta_{0},\phi_{0}):=\bigcup_{l=0}^{\infty}\bigcup_{m=-l}^{l}\left[C_{lm}\right]$, where $\bigcup$ denotes vector concatenation. Thus, the spherical harmonics coefficient vector can be used to uniquely represent a point $(\theta_{0},\phi_{0})$ on the sphere. In practice, we truncate the coefficient vector up to its leading $(L+1)^{2}$ dimensions, where $L$ is the maximum degree of associate Legendre polynomials:
@@
We call this vector the $(L+1)$-degree Spherical Harmonics Dirac Delta (SHDD) Representation of $(\theta_{0},\phi_{0})$ and $\mathbb{PE}_{\text{SHDD}}$ the SHDD encoder. Each SHDD representation corresponds to an approximation of the true spherical Dirac delta function $\delta_{(\theta_{0},\phi_{0})}$, whose probability density concentrates in a region surrounding $(\theta_{0},\phi_{0})$ rather than a single point, enabling differentiable comparison between Dirac delta functions. The Legendre polynomials have finer granularity as their degree $L$ increases, which makes SHDD representations, like other frequency-based location encoding methods such as Sphere2Vec (mai2023sphere2vec, ), capable of capturing multi-scale spatial information.
Since for spherical Dirac delta functions, the coefficients of the Legendre polynomials are the values of the Legendre polynomials at $(\theta_{0},\phi_{0})$ (arfken2011mathematical, ), i.e.,
$C_{lm}\equiv Y_{lm}(\theta_{0},\phi_{0}),\forall l,m,$ (6)
the encoding procedure can be reduced to a simple look-up of $Y_{lm}$ values.
4.3 SHDD KL-Divergence
Before describing our decoding algorithm we need to measure the difference between any $(L+1)$-degree SHDD representation of a spherical Dirac delta function $\delta$, and an arbitrary $\mathbb{R}^{(L+1)^{2}}$ vector corresponds to certain spherical function $F$. We propose to use the reverse KL-divergence between (the normalized) $F$ and $\delta$ as the difference measure $\mathcal{E}$.
Let $p_{(\theta,\phi)}$ and $q_{\mathbf{e}}$ be the normalized probability distributions corresponding to the SHDD representation of $(\theta,\phi)$ and an arbitrary $\mathbb{R}^{(L+1)^{2}}$ vector $\mathbf{e}=\bigcup_{l=0}^{L}\bigcup_{m=-l}^{l}\left[e_{lm}\right]$:
$p_{(\theta,\phi)}(u,v):=$ $\exp\left(\sum_{l=0}^{L}\sum_{m=-l}^{l}Y_{lm}(\theta,\phi)Y_{lm}(u,v)\right)/Z(\mathbb{PE}_{\text{SHDD}}(\theta,\phi))$ (7)
@@ 4.3 SHDD KL-Divergence
Here $Z(\mathbf{e})$ is a normalization constant $Z(\mathbf{e})=\int_{0}^{\pi}\int_{0}^{2\pi}\exp\left(\sum_{l=0}^{L}\sum_{m=-l}^{l}e_{lm}Y_{lm}(u^{{}^{\prime}},v^{{}^{\prime}})\right)\mathbf{d}u^{{}^{\prime}}\mathbf{d}v^{{}^{\prime}}$. Then the SHDD KL-divergence is computed by
$\mathcal{L}_{\text{SHDD-KL}}$ $(\mathbf{e},\mathbb{PE}_{\text{SHDD}}((\theta,\phi))):=\int_{0}^{\pi}\int_{0}^{2\pi}q_{\mathbf{e}}(u,v)\log\dfrac{q_{\mathbf{e}}(u,v)}{p_{(\theta,\phi)}(u,v)}\mathbf{d}u\mathbf{d}v$ (9)
It is easy to verify that the SHDD KL-divergence is a continuous difference measure. As for the property described in Equation 4, notice that by gibbs2002choosing , the Wasserstein-2 distance $W_{2}$ between $p_{(\theta,\phi)}$ and $q_{\mathbf{e}}$ is bounded by the KL-divergence in the following inequality:
@@
$C$ is a finite constant. $W_{2}$, being the Earth Mover’s Distance, quantifies the amount of probability mass transport between two distributions. Thus, when $\mathcal{L}_{\text{SHDD-KL}}(\mathbf{e},\mathbb{PE}_{\text{SHDD}}((\theta,\phi))$ is small, the difference in probability mass distribution is also small, and consequently the largest-mass-region found by the mode-seeking SHDD decoder will also remain mostly unchanged. Figure 1 visualizes this with concrete examples (pretrained Sphere2Vec location encoder and learned neural decoder v.s. our SHDD encoder and decoder) using heatmaps.
@@ 4.4 SHDD Decoding
It is impractical to compute $\mathbb{PD}_{\text{KL}}$ exactly. Luckily, Equation 6 makes a natural simplification possible. Notice that minimizing reverse KL-divergence leads to mode-seeking behavior (minka2005divergence, ), i.e. the $(\theta,\phi)$ that satisfies Equation 11 should fall within the region with the largest probability mass. Thus, we can decode $\mathbf{e}$ by finding the center of its probability mass concentration.
@@
$\mathbb{PD}_{\text{mode}}(\mathbf{e};\rho):=\operatorname*{arg\,max}_{(\theta,\phi)}\bigg\{\int_{\theta-\rho}^{\theta+\rho}\int_{\phi-\rho}^{\phi+\rho}\exp\left(\sum_{l=0}^{L}\sum_{m=-l}^{l}e_{lm}Y_{lm}(u,v)\right)\mathbf{d}u\mathbf{d}v\bigg\}$ (12)
@@
One advantage of adopting the SHDD decoder is its learning-free property. Unlike learned neural decoders, there is no loss introduced during the decoding stage. Besides, the mapping from diffusion outputs to spherical coordinates is shown to be continuous and relatively smooth. Therefore, it is safe to train latent diffusion models only using the SHDD KL-divergence loss $\mathcal{L}_{\text{SHDD-KL}}$. Another critical advantage is that the SHDD decoder does not depend on specific partitions of the spherical surface or the spatial distributions of image/location galleries. This is because the SHDD representation is in effect a continuous spherical function and in theory one can evaluate it in arbitrary resolution. The only two constraints are the maximum degree of Legendre polynomials $L$ which limits the spatial resolution of the spherical function itself and the computational resources (e.g., float32 or float64, evaluation granularity $\rho$), both being independent from other factors.
@@ 4.5 Conditional SirenNet-Based UNet (CS-UNet)
Figure 2: (a): The architecture of Condition SirenNet Module (C-Siren). $x$ is the input latent vector, $x^{{}^{\prime}}$ is the output latent vector, $t$ is the scalar timestep, and $e_{I}$ is the embedding of the input image. $d_{i}$ is the input dimension, $d_{o}$ is the output dimension, $d_{T}$ is the time embedding dimension, $d_{I}$ is the conditional embedding dimension. (b): The architecture of Conditional SirenNet-Based UNet (CS-UNet) and the workflow of LocDiff. $d$ is the latent dimension. The numbered circles denote the order of training steps.
@@
Inspired by russwurm2024locationencoding , we used SirenNet (sitzmann2020implicit, ) as the backbone of our diffusion model. The theoretical motivation behind this decision is that Spherical Harmonics coefficients are sums of sinusoidal and cosinusoidal functions (See Appendix A.3). Using sine as the activation function helps preserve gradients because the derivatives of sinusoidal/cosinusoidal functions are still sinusoidal and cosinusoidal functions. Figure 2 depicts the network architecture of the Conditional SirenNet (C-Siren) module. The design is straightforward: inputs are the latent vector $x$, the image condition embedding $e_{I}$, and the diffusion step $t$. First, we use feed-forward layers to project $x$ and $e_{I}$ into hidden vectors $h_{x}$, $h_{I}$. Then we use the sinusoidal embedding layer (song2021denoising, ) and feed-forward layers to project the discrete diffusion timestep $t$ into a scale vector $\alpha_{t}$ and a shift vector $\epsilon_{t}$. Then, we transform $h_{x}$ into $h_{x}=(1+\alpha_{t})\odot h_{x}+\epsilon_{t}$, which is an unconditional denoising step. Following that, we sum the transformed $h_{x}$ and the condition $h_{I}$ and pass the sum to a feed-forward layer, which adjusts denoising step under the guidance of the condition. Finally, output the sine-activated hidden vector to the next C-Siren module. Figure 2 describes the architecture of CS-Unet.
4.6 LocDiff
@@
Next, we introduce the training cycle of our LocDiff model as illustrated in Figure 2. A training data sample includes an input image $\mathbf{I}$ and its associated geolocation $\mathbf{p}=(\theta,\phi)$ serving as the prediction target. First, we use a frozen CLIP-based image encoder (radford2021learning, ) to encode the image $\mathbf{I}$ into an image embedding $e_{I}$. Then, we encode the geolocation $\mathbf{p}=(\theta,\phi)$ into its SHDD representation $\mathbb{PE}_{\text{SHDD}}(\theta,\phi)$ and store them in a look-up table. Following that, we perform a standard DDPM training (NEURIPS2020_4c5bcfec, ) based on the proposed CS-UNet architecture as shown in Figure 2. In a forward pass of the latent diffusion process, the spherical Dirac delta function $\delta_{(\theta_{0},\phi_{0})}$ defined by $\mathbb{PE}_{\text{SHDD}}(\theta,\phi)$ will be gradually added noise until being reduced to a vector whose values in each dimension are purely generated from Gaussian noise. In a backward pass of the latent diffusion model, the CS-UNet will start with a noise vector and gradually recover $\delta_{(\theta_{0},\phi_{0})}$. We implement the DDPM algorithm based on the open-source PyTorch implementation. We use the SHDD KL-divergence $\mathcal{L}_{\text{SHDD-KL}}$ between the ground-truth SHDD representation $\mathbb{PE}_{\text{SHDD}}(\theta,\phi)$ and the diffusion output as the training objective, because it is more computationally stable and preserves the spatial multi-scalability than the spherical MSE (e.g. great circle distance) loss. During inferencing, we sample coefficient vectors from Gaussian noise conditioned on CLIP-based image embeddings and use $\mathbb{PD}_{\text{mode}}$ to predict locations. There are two important implementation details worth mentioning. In practice, the integrals in Equation 9 and Equation 12 are approximated by summation. More specifically, we select a set of $N$ anchor points $\mathcal{A}_{N}=\{(\theta_{i},\phi_{i})\in\mathcal{C}\}_{i=1}^{N}$ on the sphere.
$\mathcal{L}_{\text{SHDD-KL}}(\mathbf{e},\mathbb{PE}_{\text{SHDD}}(\theta,\phi))=\sum_{i=1}^{N}q_{\mathbf{e}}(\theta_{i},\phi_{i})\log\dfrac{q_{\mathbf{e}}(\theta_{i},\phi_{i})}{p_{(\theta,\phi)}(\theta_{i},\phi_{i})}$ (13)
@@
$\mathbb{PD}_{\text{mode}}(\mathbf{e};\rho)=\operatorname*{arg\,max}_{(\theta,\phi)}\sum_{i=1}^{N}\mathbb{I}\{d_{\mathcal{C}}((\theta,\phi),(\theta_{i},\phi_{i}))<\rho\}\exp\left(\sum_{l=0}^{L}\sum_{m=-l}^{l}e_{lm}Y_{lm}(\theta_{i},\phi_{i})\right)$ (14)
@@
$\mathcal{L}_{\text{SHDD-KL}}$ is used for training, thus we random sample $N=2048$ anchor points over the globe for each mini-batch to avoid overfitting. As for $\mathbb{PD}_{\text{mode}}$, the choice of $\mathcal{A}_{N}$ introduces inductive bias – the regions with more anchor points may have heavier impact on the decoding results and higher spatial resolutions. However, Table 3 shows that LocDiff performs stably well on different $\mathcal{A}_{N}$. See Appendix A.5 for the pseudo codes of LocDiff.